a photon of violet light has a wavelength of 423 nm, calculate the frequency

what is the energy in kilojoules per mole?
Answer #1

Escquency & vocc - speed of Light V = 3 x 108 m/s 423 nm Im= 100 m Inma 10-9 m. 1= 3x10 m 423810-9 m V=3x10 6 +9 = 7.092 x 10